# Monte Carlo sampling with integrator snippets

### v1

@@ 9            $z_{k}=\psi_{2}(z_{k-1})$
3.5 Numerical illustration: orthant probabilities
@@
Since we are trying to sample from the constrained Gaussian distributions, the Hamiltonian equation can be solved exactly and $w_{n,k}$ is always $1$. As a result, the incremental weights for the trajectories simplify to $\Phi(\mathcal{B}_{n}(u_{n-1}))$ and each particle on the trajectory starting from $z_{t}$ will have an incremental weight proportional to $\Phi(\mathcal{B}_{n}(u_{n-1}))$. To obtain a trajectory, we follow [30] who perform HMC with reflections to sample. As the dimension increases, the number of reflections performed under $\psi_{n}$ will also increase given a fixed integration time. We adaptively tuned the integrating time $\varepsilon$ to ensure that the everage number of reflections at each SMC step does not exceed a given threshold. In our experiment we set this threshold to be $5$. To show that the waste-recycling RSMC algorithm scales well in high dimension, we set $d=150$, $a=(1.5,1.5,...)$ and $b=(\infty,\infty,...)$. Also, we use the same covariance matrix in [14] and perform variable re-ordering as suggested in [32] before the simulation.
@@
Figures 9 and 10 show the results obtained with $N\times(T+1)=50,000$ and various values for $N$. With a quarter of the number of particles used in [14], the waste-recycling HSMC algorithm achieves comparable performance when estimating the normalising constant (i.e. the orthant probability). Moreover, the estimates are stable for different choices of $N$ values, although one observes that the algorithm achieves best performance when $N=500$ (i.e. each trajectory contains $100$ particles). This also suggests that the integrating time should be adaptively tuned in a different way to achieve the best performance given a fixed computational budget. Estimates of the function $\varphi(x_{0:d})=\mathbb{E}(1/d\sum_{i=1}^{d}x_{i})$ with respect to the Gaussian distribution $\mathcal{N}_{d}(\mathbf{0},\Sigma)$ truncated between $\mathbf{a}$ and $\mathbf{b}$ are also stable for different choices of $N$, although they are more variable than those obtained in [14]. This indicates that the waste-recycling HSMC algorithm does scale well in high dimension. We note that this higher variance compared to the waste-free SMC of [14], is obtained in a scenario where they are able to exploit the particular structure of the problem and implement an exact Gibbs sampler to move the particles. The integrator snippet SMC we propose is however more general and applicable to scenarios where such a structure is not present.
4 Preliminary theoretical exploration
@@
4.1 Variance using folded mixture samples
This section establishes variance reduction for an estimator of $\mu(f)$ of the type (15), but where it is assumed that $\{\check{z}_{n}^{(i)},i\in\llbracket N\rrbracket\}$ are iid distributed according to $\bar{\mu}_{n}$. While this is not the exact distribution in the algorithm this provides a proxy representative of the quantities one needs to control when analysing SMC samplers [15, Chapter 9]: variance of estimators in an SMC can be decomposed as the sum of local variances at each iterations, which convergence to the variance terms considered hereafter as $N\rightarrow\infty$. The main message is that $\{f\circ\psi^{k}(z),k\in\llbracket P\rrbracket\}$ can be understood as playing the rôle of control variates, with the potential of reducing variance.
We use the following simplified notation throughout. Let
$\bar{\mu}(k,{\rm d}z):=\frac{1}{T+1}\mu_{k}({\rm d}z)\,,$
with $\mu_{k}:=\mu^{\psi^{-k}}$. The fundamental property exploited throughout is (11), which can be rephrased as, for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu$-integrable,
$\int\Bigl{\{}\sum_{k=0}^{T}f\circ\psi^{k}(z)\bar{\mu}(k\mid z)\Bigr{\}}\bar{\mu}({\rm d}z)=\mu(f)\,,$
or written differently with $\bar{f}(k,z):=f\circ\psi^{k}(z)$
$\mathbb{E}_{\bar{\mu}}\left(\mathbb{E}_{\bar{\mu}}(\bar{f}(\mathrm{T},\check{Z})\mid\check{Z})\right)=\mathbb{E}_{\bar{\mu}}\left(\bar{f}(\mathrm{T},\check{Z})\right)=\mathbb{E}_{\mu}\left(f(Z)\right).$
The estimator we use is therefore a so-called “Rao-Blackwellized estimator”, assuming $\check{Z}^{(1)},\ldots,\check{Z}^{(N)}\overset{{\rm iid}}{\sim}\bar{\mu}$,
of variance ${\rm var}_{\bar{\mu}}\left(\mathbb{E}_{\bar{\mu}}\big{(}\bar{f}(\mathrm{T},\check{Z})\mid\check{Z}\big{)}\right)/N$. The following relates this variance to that of the standard estimator using $N$ iid samples from $\mu$, which would likewise play a role in the asymptotic properties of SMC algorithms.
Proposition 15.
We have
${\rm var}_{\bar{\mu}}\left(\mathbb{E}_{\bar{\mu}}\big{(}\bar{f}(\mathrm{T},\check{Z})\mid\check{Z}\big{)}\right)={\rm var}_{\mu}\left(f(\check{Z})\right)-\mathbb{E}_{\bar{\mu}}\big{(}{\rm var}_{\bar{\mu}}(\bar{f}(\mathrm{T},\check{Z})\mid\check{Z})\big{)}$
Proof.
The variance decomposition identity yields
but from the fundamental property
$=\mathbb{E}_{\mu}\left(f(\check{Z})^{2}\right)-\mathbb{E}_{\mu}\left(f(\check{Z})\right)^{2}$
$={\rm var}_{\mu}\left(f(\check{Z})\right).$
The following provides us with a notion of effective sample size for estimators of the form $\check{\mu}(f)$.

### v2

@@ 9            $z_{k}=\psi_{2}(z_{k-1})$
Appendix E Early experimental results: HMC-IS
E.1 Numerical illustration: logistic regression
$L(x)=\prod_{i}^{n}F(z_{i}^{\top}x\cdot y_{i}),$ (38)
For both algorithms, we set the total number of particles at each SMC step to be $N(T+1)=10,000$. For the waste-free SMC, the Markov kernel is chosen to be a random-walk Metropolis-Hastings kernel with covariances adaptively computed as $2.38^{2}/d\ \hat{\Sigma}$, where $\hat{\Sigma}$ is the empirical covariance matrix obtained from the particles in the previous SMC step. For the Hamiltonian Snippet-SMC algorithm, we set $\psi_{n}$ to be the one-step leap-frog integrator with stepsize $\epsilon$, $U_{n}(x)=-\log(\pi_{n}(x))$ and $\varpi$ a $\mathcal{N}(0,\mathrm{Id})$. To investigate the stability of our algorithm, we ran Hamiltonian Snippet SMC with $\epsilon=0.05,0.1,0.2$ and $0.3$. For both algorithms, the temperatures $\lambda_{n}$ are adaptively chosen so that the effective sample size (ESS) of the current SMC step will be $\alpha ESS_{max}$, where $ESS_{max}$ is the maximum ESS achievable at the current step. In our experiments, we have chosen $\alpha=0.3,0.5$ and $0.7$ for both algorithms.
Performance comparison
Figure 23 shows the boxplots of estimates of the logarithm of the normalising constant obtained from both algorithms, for different choices of $N$ and $\epsilon$ for the Hamiltonian Snippet SMC algorithm. The boxplots are obtained by running both algorithms 100 times for different of the algorithm parameters, with $\alpha=0.5$ in all setups. Several points are worth observing. For a suitably choice of $\epsilon$, the Hamiltonian Snippet SMC algorithm can produce stable and consistent estimates of the normalising constant with $10,000$ particles at each iteration. On the other hand, however, the waste-free SMC algorithm fails to produce accurate results for the same computational budget. It is also clear that with larger values of $N$ (meaning smaller value of $T$ and hence shorter snippets), the waste-free SMC algorithm produces results with larger biases and variability. For Hamiltonian Snippet SMC algorithm, the results are stable both for short and long snippets when $\epsilon$ is equal to $0.1$ or $0.2$. Another point is that when $\epsilon=0.05$ or $0.3$, the Hamiltonian Snippet SMC algorithm becomes unstable with short (i.e. $\epsilon=0.05$) and long (i.e. $\epsilon=0.3$). Possible reasons are for too small a stepsize the algorithm is not able to explore the target distribution efficiently, resulting in unstable performances. On the other hand, when the stepsize is too large, the leapfrog integrator becomes unstable, therefore affecting the variability of the weights; this is a common problem of HMC algorithms. Hence, a long trajectory will result in detoriate estimations. Hence, to obtain the best performance, one should find a way of tuning the stepsize and trajectory length along with the SMC steps.
Figure 23: Estimates of the normalising constant (in log scale) obtained from both Hamiltonian Snippet SMC and waste-free SMC algorithm. Left: Estimate obtained from Hamiltonian Snippet SMC algorithm with different values of $\epsilon$. Right: Estimates obtained from waste-free SMC algorithm.
Figure 24: Estimates of the mean of marginals obtained from both Hamiltonian Snippet SMC and the waste-free SMC algorithms. Left: Estimate obtained from the Hamiltonian Snippet SMC algorithm with difference values of $\epsilon$. Right: Estimates obtained from the waste-free SMC algorithm.
Computational Cost
Figure 25: Simulation times for both algorithms. Left: Simulation time for Hamiltonian Snippet SMC algorithm, with $N=100,\epsilon=0.2,\alpha=0.5$. Right:Simulation times for the waste-free SMC with $N=100$
E.2 Numerical illustration: orthant probabilities
@@
Since we are trying to sample from the constrained Gaussian distributions, the Hamiltonian equation can be solved exactly and $w_{n,k}$ is always $1$. As a result, the incremental weights for the trajectories simplify to $\Phi(\mathcal{B}_{n}(u_{n-1}))$ and each particle on the trajectory starting from $z_{t}$ will have an incremental weight proportional to $\Phi(\mathcal{B}_{n}(u_{n-1}))$. To obtain a trajectory, we follow [51] who perform HMC with reflections to sample. As the dimension increases, the number of reflections performed under $\psi_{n}$ will also increase given a fixed integration time. We adaptively tuned the integrating time $\epsilon$ to ensure that the everage number of reflections at each SMC step does not exceed a given threshold. In our experiment we set this threshold to be $5$. To show that the waste-recycling RSMC algorithm scales well in high dimension, we set $d=150$, $a=(1.5,1.5,...)$ and $b=(\infty,\infty,...)$. Also, we use the same covariance matrix in [26] and perform variable re-ordering as suggested in [54] before the simulation.
@@
Figures 26 and 27 show the results obtained with $N\times(T+1)=50,000$ and various values for $N$. With a quarter of the number of particles used in [26], the waste-recycling HSMC algorithm achieves comparable performance when estimating the normalising constant (i.e. the orthant probability). Moreover, the estimates are stable for different choices of $N$ values, although one observes that the algorithm achieves best performance when $N=500$ (i.e. each trajectory contains $100$ particles). This also suggests that the integrating time should be adaptively tuned in a different way to achieve the best performance given a fixed computational budget. Estimates of the function $\varphi(x_{0:d})=\mathbb{E}(1/d\sum_{i=1}^{d}x_{i})$ with respect to the Gaussian distribution $\mathcal{N}_{d}(\mathbf{0},\Sigma)$ truncated between $\mathbf{a}$ and $\mathbf{b}$ are also stable for different choices of $N$, although they are more variable than those obtained in [26]. This indicates that the waste-recycling HSMC algorithm does scale well in high dimension. We note that this higher variance compared to the waste-free SMC of [26], is obtained in a scenario where they are able to exploit the particular structure of the problem and implement an exact Gibbs sampler to move the particles. The integrator snippet SMC we propose is however more general and applicable to scenarios where such a structure is not present.
Appendix F Early experimental results: filamentary distributions
F.1 Numerical illustration: simulating from filamentary distributions
:=$\mathcal{M}:=\left\{x\in\mathsf{X}:\ell(x)=0\right\},$
$\pi_{\epsilon}(x)\propto k_{\epsilon}\circ\ell(x)\,\pi(x)\,.$
[…]
Integrator Snippet SMC
[…]
In Fig. 32 we illustrate robustness of integrator snippet SMC to the choice of $\epsilon$ and $T$ through the support covered; note that the computational cost of the three algorithms considered is comparable.
Figure 32: Scatter plot of the first two coordinates of $\{x_{n,k}^{(i)},i\in\llbracket N\rrbracket,k\in\llbracket 0,T\rrbracket\}$ for three SMC algorithms: Snippet SMC (traj), Snippet SMC using the snippets endpoints only (endp) and the SMC sampler (smc). The stepsize is decreased from a high value for rows 1-3 $(\epsilon=1.0)$, to lower values for rows 4-6 ($\epsilon=0.1$) and rows 7-9 $(\epsilon=0.01)$. The three columns correspond to different snippet length, $T=10,30,50$.
F.2 Going round the bend: bananas and pears
In this section we illustrate how ODEs and associated integrators can be used in order to update states involved in simulation algorithm for probability distributions involving geometrically awkward densities. The banana shaped distribution introduced by [34] in the context of adaptive MCMC algorithm is a prototypical toy example of dependencies encountered in some practical situation e.g. inverse problems. The banana shaped distribution defined on $\mathsf{X}=\mathbb{R}^{d},$ has density with respect to the Lebesgue measure $\pi(x)\propto\mathcal{N}\big{(}\phi(x);0,\text{$\Sigma$}\big{)}$ where $\phi(x)=(x_{1},x_{2}+0.03(x_{1}^{2}-100),x_{2},\ldots,x_{n})$ and $\Sigma={\rm diag}(100,1,\ldots,1)$ – see its contours for $d=2$ in Fig. 33. A first suggestion could be to follow the contours of $\pi$ using the update described in the previous section. Here we suggest the following flow, for $v\in\{-1,1\}$,
$t$$\mapsto\begin{cases}x_{1}(t)=vt\\ x_{2}(t)=-0.03(x_{1}^{2}(t)-100)+x_{2}(0)+0.03(x_{1}^{2}(0)-100)\\ x_{i}(t)=x_{i}(0)&\text{for }i>2\end{cases}$
Motivation for this is that $t\mapsto x_{2}(t)+0.03(x_{1}^{2}(t)-100)=x_{2}(0)+0.03(x_{1}^{2}(0)-100)$ is a constant, effectively unbending the banana and effectively applying a standard Gibbs sampler [64] on a Gaussian in a different set of coordinates. This suggests a practical implementation of generalized forms of the Gibbs sampler, which can be traced back to [18] and were recently revived in [52]. Naturally in practice such updates must be completed with other updates to ensure irreducibility.
In more general situations where $\phi(x)=(x_{1},u(x_{1},x_{2}),x_{3},\ldots)$ for some $u\colon\mathsf{X}^{2}\rightarrow\mathbb{R}$ such that $\pi$ is indeed a probability density, the idea can be generalized with the ideal aim of keeping $t\mapsto u\big{(}x_{1}(t),x_{2}(t)\big{)}$ or $t\mapsto x_{1}(t)^{2}+u\big{(}x_{1}(t),x_{2}(t)\big{)}^{2}$ constant, either using THUG or even Hamilton’s equation as suggested by the last contraint. Note however that non-separability may cause additional numerical difficulties in the latter scenario.
Figure 33: The banana distribution, $d=2$
Another popular illustrative toy example is Neal’s funnel on which numerous standard algorithms may struggle to converge. The funnel consists of a mixture of normal distributions
$\pi(x,y)=\mathcal{N}(y;0,3)\mathcal{N}(x;0,\exp(y/2))\,,$
defining a probability density on $\mathbb{R}^{2}$ of contours shown in Fig. 34–multiple $x$’s can be incorporated therefore taking the form of a hierchical model. It exemplifies some of the difficulties which may be encountered when sampling from hierarchical models.
Figure 34: Neal’s funnel
The narrow funnel is the cause of failure of traditional algorithms. A simple idea here consists of identifying a flow such that $t\mapsto\mathcal{N}(x_{t};0,\exp(y_{t}/2))\equiv\mathcal{N}(x_{0};0,\exp(y_{0}/2))$ for given $(x_{0},y_{0})$. A straightforward choice imposes $\dot{y}_{t}=v$ with $v\in\{-1,1\}$ and $x_{t}/\exp(y_{t}/4)=x_{0}/\exp(y_{0}/4)$ for $t>0$, therefore ensuring that $x_{t}\sim\mathcal{N}(0,\exp(y_{t}/2))$ for $t>0$. This can be solved analytically
$x_{t}$ $=x_{0}\exp\left(y_{0}/4+y_{t}/4\right)=x_{0}\exp\left((y_{0}+vt)/4\right)$
$y_{t}$ $=vt\,.$
This corresponds to a simple reparametrizations, standard in the literature. However our suggested use differs in that both components are updated simultaneously within an MCMC kernel, and illustrate the potential of this type of ideas. Such a dynamic should naturally be mixed with other updates ensuring irreducibility.
Appendix G Adaptation: implementational details
G.1 About the structure of the mother distribution
It is useful to provide details about the (hierchical) structure of the target distributions useful for implementational purposes. The mother distribution in (20) is given by
$\bar{\mu}\big{(}k,{\rm d}(\epsilon,z);\gamma,\theta\big{)}:=\frac{1}{T+1}\mu^{\psi_{\epsilon}^{-k}}({\rm d}z;\gamma)\nu_{\theta}({\rm d}\epsilon)$
from which we notice that
$\bar{\mu}\big{(}k,{\rm d}(\epsilon,z);\theta,\gamma\big{)}$ $=\bar{\mu}\big{(}k,{\rm d}z\mid\epsilon,\gamma\big{)}\nu_{\theta}({\rm d}\epsilon)$
$=\bar{\mu}\big{(}k\mid\epsilon,z;\gamma\big{)}\bar{\mu}\big{(}{\rm d}z\mid\epsilon;\gamma\big{)}\nu_{\theta}({\rm d}\epsilon)$
that is the first conditionals do not depend on $\theta\in\Theta$. Indeed, marginalising yields $\bar{\mu}\big{(}{\rm d}\epsilon,z;\theta,\gamma\big{)}=\nu_{\theta}({\rm d}\epsilon)$, implying
$\bar{\mu}\big{(}k,{\rm d}z\mid\epsilon;\gamma,\theta\big{)}$ $=\frac{1}{T+1}\mu^{\psi_{\epsilon}^{-k}}({\rm d}z;\gamma)$
and justifying the notational simplifications, leading to
$\bar{\mu}\big{(}{\rm d}z\mid\epsilon;\gamma,\theta\big{)}=\bar{\mu}\big{(}{\rm d}z\mid\epsilon;\gamma\big{)}=\frac{1}{T+1}$ $\sum_{l=0}^{T}\mu^{\psi_{\epsilon}^{-l}}({\rm d}z;\gamma)$
$\bar{\mu}\big{(}k\mid\epsilon,z;\gamma,\theta\big{)}=\bar{\mu}\big{(}k\mid\epsilon,z;\gamma\big{)}=$ $\frac{{\rm d}\mu^{\psi_{\epsilon}^{-k}}}{{\rm d}\big{(}\sum_{l=0}^{T}\mu^{\psi_{\epsilon}^{-l}}\big{)}}(z;\gamma)\,.$
With $\bar{f}=\bar{f}(\epsilon,K,Z)$ for $\bar{f}(\epsilon,k,z):=f\circ\psi_{\epsilon}^{k}(z)$ for $f\colon\mathsf{Z}\rightarrow\mathbb{R}^{m}$ for $m\in\mathbb{N}$,
$\bar{\mu}\big{(}\bar{f}\mid\epsilon,z;\gamma\big{)}:=\sum_{k=0}^{T}\bar{\mu}\big{(}k\mid\epsilon,z;\gamma\big{)}f\circ\psi_{\epsilon}^{k}(z)\,.$
G.2 Kullback-Leibler minimisation
To minimize the KL divergence to determine $\theta_{n}$ we first note that, with $C$ independent of $\theta$ and possibly changing on each appearance,
$\theta\mapsto{\rm{\rm KL}}\big{(}\tilde{\mu}(\cdot;\gamma_{n},\theta_{n}),\bar{\mu}(\cdot\mid\gamma_{n},\theta)\big{)}$
$=C+\int\log$ $\left(\frac{v_{\gamma_{n-1}}(\epsilon,z)\nu_{\theta_{n-1}}({\rm d}\epsilon)\bar{\mu}\big{(}{\rm d}z\mid\epsilon;\gamma_{n1}\big{)}}{\nu_{\theta}({\rm d}\epsilon)\bar{\mu}\big{(}{\rm d}z\mid\epsilon;\gamma_{n}\big{)}}\right)\tilde{\mu}\big{(}{\rm d}(\epsilon,z);\gamma_{n},\theta_{n}\big{)}$
$=C-\int\log$ $\left(\nu_{\theta}(\epsilon)\right)\tilde{\mu}\big{(}{\rm d}(\epsilon,z);\gamma_{n},\theta_{n}\big{)}$
$=C-\int\log$ $\left(\nu_{\theta}(\epsilon)\right)\,\tilde{\mu}\big{(}{\rm d}\epsilon;\gamma_{n},\theta_{n}\big{)}\,.$
We remark that minimization of the above is easily achieved in scenarios where $\nu_{\theta}$ is chosen in the exponential family, e.g. $\nu_{\theta}$ has density of the form $\nu_{\theta}(\epsilon)=\exp\big{(}\langle\theta,T(\epsilon)\rangle-A(\theta)\big{)}$, in which case minimisation of (21) amounts to solving
$\theta\mapsto\mathbb{E}_{\nu_{\theta}}\left(T(\epsilon)\right)=\mathbb{E}_{\tilde{\mu}\big{(}\cdot;\gamma_{n},\theta_{n}\big{)}}\left(T(\epsilon)\right)\,.$
In practice the expectation with respect to $\tilde{\mu}\big{(}\cdot\gamma_{n},\theta_{n}\big{)}$ is obtained using self-normalised importance sampling of the seed particles $\{(\check{z}_{n}^{(i)},\check{\epsilon}_{n}^{(i)}),i\in\llbracket N\rrbracket\}$ representing $\bar{\mu}_{n}$ or weighted samples $\{(z_{n-1}^{(i)},\epsilon_{n-1}^{(i)}),i\in\llbracket N\rrbracket\}$ (see Alg. (3) and surrounding comments), leading in the exponential family scenario to solving
$\theta\mapsto\mathbb{E}_{\nu_{\theta}}\left(T(\epsilon)\right)=\frac{\sum_{i=1}^{N}T(\check{\epsilon}_{n}^{(i)})v(\check{\epsilon}_{n}^{(i)}),\check{z}_{n}^{(i)})}{\sum_{i=1}^{N}v(\check{\epsilon}_{n}^{(i)},\check{z}_{n}^{(i)})}\,.$
In the general scenario minimization is equivalent to maximum likehood estimation with samples $\{(\check{z}_{n}^{(i)},\check{\epsilon}_{n}^{(i)}),i\in\llbracket N\rrbracket\}$ as data.
Now we derive a more expression for $v_{\gamma}(\epsilon,z):=\mathrm{Tr}\big{[}{\rm var}\big{(}\bar{f}\mid\epsilon,z;\gamma\big{)}\big{]}$ useful for implementation.
$v_{\gamma}(\epsilon,z)$ $=\mathrm{Tr}\{\mathbb{E}\big{[}\bar{(f}-\mathbb{E}(\bar{f}\mid\epsilon,z))\bar{(f}-\mathbb{E}(\bar{f}\mid\epsilon,z))^{\top}\mid\epsilon,z;\gamma\big{]}\}$
$=\mathbb{E}\{{\rm Tr}\big{[}\bar{(f}-\mathbb{E}(\bar{f}\mid\epsilon,z))\bar{(f}-\mathbb{E}(\bar{f}\mid\epsilon,z))^{\top}\big{]}\mid\epsilon,z;\gamma\}$
$=\mathbb{E}\{\|\bar{f}-\mathbb{E}(\bar{f}\mid\epsilon,z)\|^{2}\mid\epsilon,z;\gamma\}$
$=\frac{1}{T+1}\sum_{k=0}^{T}\Bigl{\|}f\circ\psi_{\epsilon}^{k}(z)-\bar{\mu}(f\mid\epsilon,z;\gamma,\theta)\Bigr{\|}^{2}\bar{\mu}\big{(}k\mid\epsilon,z;\gamma,\theta\big{)}\,.$ (43)
G.3 Calculations in the inverse Gaussian scenario
The Inverse Gaussian with mean $\theta>0$ and shape parameter $\lambda>0$ has skewness $s=3\sqrt{\theta/\lambda}$. Therefore for a fixed skewness $s>0$ and mean parameter $\theta>0$ the IG distribution has density
$\nu_{\theta}(\epsilon;s)=\sqrt{\frac{9\theta}{2\pi\epsilon^{3}s^{2}}}\exp\left(-\frac{9\theta(\epsilon-\theta)^{2}}{2\theta^{2}\epsilon s^{2}}\right)\,,$
w.r.t. the Lebesgue measure. We aim to minimise the Kullback-Leibler divergence ${\rm\theta\mapsto KL}(\nu,\nu_{\theta})$ for a given probability distribution. The corresponding log-likelihood is
$\log\nu_{\theta}(\epsilon;s)=\log(3)+\frac{1}{2}\log(\theta)-\frac{1}{2}\log(2\pi)-\frac{3}{2}\log(\epsilon)-\log(s)-\frac{9(\epsilon^{2}+\theta^{2}-2\epsilon\theta)}{2\theta\epsilon s^{2}}\,,$
and its derivative w.r.t. $\theta>0$ is
$\frac{{\rm d}}{{\rm d}\theta}\log\nu_{\theta}(\epsilon;s)$ $=\frac{1}{2\theta}-\frac{9}{2\epsilon s^{2}}\left[\frac{{\rm d}}{{\rm d}\theta}\left(\frac{\epsilon^{2}}{\theta}+\theta-2\epsilon\right)\right]$
$=\frac{1}{2\theta}-\frac{9}{2\epsilon s^{2}}\left[-\frac{\epsilon^{2}}{\theta^{2}}+1\right]$
$=\frac{1}{2\theta}-\frac{9(\theta^{2}-\epsilon^{2})}{2\epsilon s^{2}\theta^{2}}\,.$
Therefore
$\frac{{\rm d}}{{\rm d}\theta}{\rm KL}(\nu,\nu_{\theta})=\frac{1}{2\theta^{2}}\frac{9}{s^{2}}\left\{\frac{s^{2}}{9}\theta-\theta^{2}\mathbb{E}_{\nu}[\epsilon^{-1}]+\mathbb{E}_{\nu}[\epsilon]\right\}=0\,,$
which has unique positive solution
$\hat{\theta}=\frac{{\displaystyle\frac{s^{2}}{9}+\sqrt{\frac{s^{4}}{81}+4\mathbb{E}_{\nu}[\epsilon^{-1}]\mathbb{E}_{\nu}[\epsilon]}}}{2\mathbb{E}_{\nu}[\epsilon^{-1}]}$
Using that $\mathbb{E}_{\nu}[\epsilon^{-1}]=\mathbb{E}_{\nu}[\epsilon]^{-1}+\lambda^{-1}$, $\lambda=9\mathbb{E}_{\nu}[\epsilon]/s^{2}$ we deduce $\mathbb{E}[\epsilon^{-1}]\mathbb{E}[\epsilon]=(9+s^{2})/9$ yielding the simpler expression
$\hat{\theta}=9\mathbb{E}_{\nu}[\epsilon]\left[\frac{{\displaystyle\frac{s^{2}}{9}+\sqrt{\frac{s^{4}}{81}+4\frac{(9+s^{2})}{9}}}}{2(9+s^{2})}\right]$
$=\mathbb{E}_{\nu}[\epsilon]\frac{s^{2}+\sqrt{(s^{2}+18)^{2}}}{2(9+s^{2})}\,.$
The constrained minimisation problem is therefore akin to a moment matching problem.
Appendix H Additional theoretical properties
@@
H.1 Variance using folded mixture samples
The following provides us with a notion of effective sample size for estimators of the form $\check{\mu}(f)$ as defined in Proposition 5.